A 50 kg boy rides a 20 kg bicycle.  The boy rides up a hill so that his height increases by 120 m in 30 seconds.. This means he created ______ Watts of power.  


1
Expert's answer
2021-01-29T05:20:11-0500

Calculate how much gravitational potential energy the kid gained:


"E=(M+m)gh."

The power is


"P=\frac{E}{t}=\frac{(M+m)gh}{t}=\frac{(50+20)9.8\cdot120}{30}=2744\text{ W}."